OpenStudy (stbest):

Please Explain This To Me What is the measure of angle L to the nearest whole degree. Link Below

OpenStudy (stbest):

\(\color{Green}{PLZ~HELP}\)

OpenStudy (stbest):

I know the Equation But my calculator wont work properly\[\sin^{-1} (\frac{ 6 }{ 7 })~~or~~\sin^{-1}(1.167)\]

OpenStudy (stbest):

\(\rlap\huge\color{red}{IT~DOESN'T~WORK}\rlap{\huge\color{red}{IT~Doesn't~WORK}}\)

OpenStudy (erak):

You would not use sin, as that needs (opposite/hypotenuse). Which two sides do you have?

OpenStudy (stbest):

Oppisite and Adjacent

OpenStudy (stbest):

so i use tan??

OpenStudy (erak):

Yes

OpenStudy (stbest):

\[\tan^{-1} (1.167)\]

OpenStudy (stbest):

which is 59.08 or rounded to 59

OpenStudy (stbest):

oh nvm

OpenStudy (stbest):

49 i mean i didnt put 1 in the calculator

OpenStudy (erak):

Just stick with arctan (7/6), no need to round, but they should give similar answers

OpenStudy (stbest):

THanks I just forgot to use which is which

OpenStudy (stbest):

\(\Huge\color{Cyan}{ERAK}\)

OpenStudy (stbest):

\(\Huge\color{Green}{Thanks}\)